OpenStudy (anonymous):

Rewrite with only sin x and cos x. cos 2x - sin x helppp

OpenStudy (anonymous):

well cos (2x) = cos^2(x) - sin^2(x) and cos^2(x) + sin^2(x) = 1 so that means cos^2(x) = 1 - 2 sin^2(x) so you could rewrite the original as 1 - 2 sin^2(x) - sin(x) does this help?

OpenStudy (anonymous):

cos2x - sin2x - sin x cos2x - sin3x cos2x + sin2x + sin x cos2x - 3 sin x these are my choices. i feel like thats the simplified form of one of these im just not sure how to simplify them.

OpenStudy (anonymous):

hm.. is the original expression \[\cos(2x) - \sin (x)\] ?

OpenStudy (anonymous):

yes